OpenStudy (anonymous):

radical of 20a^2b Simplify by factoring.

OpenStudy (jlastino):

2a sqrt 5b

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt{20a^2b}\] \[\Large \sqrt{4*5a^2b}\] \[\Large \sqrt{4}\sqrt{5}\sqrt{a^2}\sqrt{b}\] \[\Large 2\sqrt{5}\times a\sqrt{b}\] \[\Large 2a\sqrt{5b}\] So \[\Large \sqrt{20a^2b}=2a\sqrt{5b}\]